Circular No. 7566

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001L IN MCG –01-30-11 A. Friedman, W. D. Li, and R. Chornock, University of California at

Berkeley, report the discovery by LOTOSS (cf. IAUC 7514) of a supernova (mag ∼ 15.4) on an unfiltered image taken on Jan. 18.4 UT (and confirmed at mag ∼ 15.7 on an earlier image taken on 2000 Dec. 31.4) with the 0.8m Katzman Automatic Imaging Telescope. SN 2001L is located at α = 11^{h}36^{m}48^{s}.83, δ = −8^{o}35′07′′.0 (equinox 2000.0), which is 0′′.3 east and 3′′.8 north of the nucleus of MCG –01-30-11. A KAIT image taken on 2000 Dec. 23.4 showed nothing at this position (limiting mag ∼ 19.0).

S. Jha, T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum of SN 2001L, obtained by M. Calkins on Jan. 19.46 UT with the F. L. Whipple Observatory 1.5m telescope (+ FAST spectrograph), identifies it as a type-Ia supernova ∼ 1 week after maximum brightness. With the NASA/IPAC Extragalactic Database recession velocity for the host galaxy of 4567 km/s, the expansion velocity of the supernova is 9400 km/s, measured using the Si II (rest 635.5 nm) feature. The spectral-feature age (Riess et al. 1997, A.J. 114, 722) of the supernova is 8 ± 2 days after maximum light.

SUPERNOVAE 2001D, 2001E, AND 2001J Jha et al. also report that a spectrum of SN 2001D (cf. IAUC 7557),

obtained as above on Jan. 19.47 UT, shows it to be a type-II supernova at or just after maximum light. The spectrum exhibits a slightly blue continuum with well-developed P-Cyg Balmer and He I lines. The expansion velocity of the supernova, measured using the Hα profile, is 10 100 km/s, where the recession velocity of IC 728 was measured to be 8500 km/s from narrow, superimposed Hα emission. A spectrum of SN 2001E (cf. IAUC 7557), taken as above on Jan. 19.49, indicates it to be a type-Ia supernova at maximum light. Superimposed, narrow Hα emission provides a recession velocity for NGC 3905 of 5790 km/s, implying an expansion velocity for the supernova of 14 200 km/s, measured via the strong Si II (rest 635.5 nm) profile. The spectral-feature age of the supernova is 0 ± 2 days after maximum light. A spectrum of SN 2001J (cf. IAUC 7564), taken as above on Jan. 19.45, shows it to be a type-II supernova at maximum light. The spectrum features a flat continuum with broad Hα, Hβ, and He I profiles. The expansion velocity based on the Hα profile is 13 000 km/s, using a recession velocity for UGC 4729 of 3860 km/s, as measured from narrow, superimposed Hα emission.

2001 January 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green